A Bibliography of Publications of John A. Nelder

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2026 Version 1.294

Abstract

This bibliography records publications of John A. Nelder.

Title word cross-reference

2^{n} [Nel63b]. ^{2} [Baş26]. H [Boo07, Dry10, LNN07, Rob07].

-Likelihood [Rob07, Boo07, Dry10, LNN07].

16 [Nel98a]. 19 [Nel01b].

20 [Nel69, Nel94c]. 2010 [Sen19]. 2nd [Che90].

3 [Col79]. 3289 [Nel01b].

8 [Nel68a].

9 [Nel68b]. 94g [Nel94c]. 96 [Nel76].

A. [Les97, LR97, Nel97, Wir97]. ab-initio [PP23]. Acceleration [YCH24]. accurate [CS03]. acoustics [KNCS21]. Adams [Ano07]. adaptive [GH12, Meh20, WYWW15]. Additive [Nel68b, Nel86a]. Afifi [Nel79b]. Algorithm [CGVP15, Gal24, Kel99, LPW12, Nel76, TFB21, DW87, YCH24, Arn93, AT18, Baş26, BPT06, FBP19, GH12, KBČ20, KNCS21, LTJ95, LL12, MM26, Meh20, MJ10, NZE^{+}15, NT02, PCB02, PB26, RWJ13, SGK16, SS04, SEABZ24, Tom95, TA13, WYWW15, WSKF13, ZZP^{+}22, Nel68b, Nel68a, Nel69, Nel76]. Algorithms [Nel68b, Nel68a, Nel69, NCC^{+}70, Nel76, Pay04, CS03, CS05, Mas06, Mas07, Mas09a, Mas09b, SGK16]. Alternative [Nel62a, Nel85b, Nel94a, LN06b]. Altman [Nel01b]. Amongst [Rip04]. Analysis [Boo07, DCZ^{+}19, Nel75b, Nel79d, Nel79c, Nel79a, Nel79b, Nel87b, NGMS94, NL95, Nel00c, Rob07, BC97, Dry10, HNX03, LN98, Nel51, Nel65a, Nel65b, Nel94c, Nel98a, Pay04, SS01, Nel72c, Nel78a, Nel72b]. Analytical [CCG12]. annealing [TA13]. Announcement [Nel75a]. ANOVA [CST04]. Application [IBV^{+}22, Nel77a, OT09, BC97, KBČ20, LN00a, MM26, Nel98a, Nel56]. approach [PP23, Nel79b]. areas [Nel93b]. Argumented [YCH24]. arising [HNX03]. Arithmetic [Aki77]. Array [Nel69]. arrays [Mas07]. ascent [Arn93]. Ashworth [Ano11, MGP11, Sen19]. Aspects [Nel93a]. August [Sen19]. Azen [Nel79b].

B [Nel79a, Nel87a]. back [Ait95, Gow95, Nel94d, Nel95a, Nel95b, RTW95, Sea95, Van95]. Baker [Alt80, Col79]. Balance [Sen04]. balanced [Nel68d]. band [DCZ^{+}19]. barrier [Pri21]. based [AT18, Baş26, DCZ^{+}19, Lee04, LL12, Meh20, NT02, WYWW15, XD14, ZZP^{+}22]. basic [Van95]. basics [Nel94d, Nel95b, Ait95, Gow95, Nel95a, RTW95, Sea95]. basis [WYWW15]. bat [SEABZ24]. bat-Nelder [SEABZ24]. Bayesian [Nel98a, BC97]. Behavioural [Nel79c]. Besag [Ano11]. between [Arn93, LN00a]. Beyond [Lee04]. bilevel [RWJ13]. Biology [Nel75e, Nel71a]. Biometry [Nel63a]. Bliss [Nel71a]. block [Nel65a, Nel65b, Nel65a]. BMDP [Nel78b]. Book [Alt80, Ano07, Boo07, Bur91, Che90, Col79, Dal72, Dry10, Gar71, Gho08, Inn70, K.71, Lac90, Nel56, Nel63a, Nel64a, Nel68c, Nel71a, Nel72b, Nel72c, Nel75b, Nel77a, Nel78a, Nel79d, Nel79c, Nel79a, Nel79b, Nel80b, Nel80a, Nel87a, Nel94b, Pla84, Rob07, Sch92, Wat86, Wil84]. Both [Sen04]. boundary [Mas09a]. bounded [WSKF13]. Brian [Nel90b]. Broemeling [Nel98a]. Building [Nel68e].

C [Dal72, Gar71, Inn70, K.71, Nel64a]. calibration [WSKF13]. Categorical [Nel75b]. centroid [FBP19]. Certain [Nel72b, HNX03]. Chambers [Nel78a]. Characteristics [CCG12]. Chatterjee [Nel79a]. Chebyshev [Meh20]. Chester [Nel71a]. cities [KNCS21]. class [Nel98c, RWJ13]. Classes [Rip04]. classical [Nel74a]. clinical [Lar97]. cliniques [Nel64a]. code [PP23]. coded [MM26]. Coefficient [Nel94b]. combination [Nel68d]. combinations [Lar97]. combining [CS05]. Comment [Car97, Nel78b, Nel86a, Wir97, Nel94c]. Comments [Ait95, Gow95, RTW95, Sea95, Nel95a, Van95]. Comparison [TFB21, Arn93]. comparisons [NL92]. Complexity [SS01]. Component [Nel71c]. components [Nel54a]. Computation [Gol69, MN14, Dal72, Gar71, Inn70, K.71]. computational [Pay04, Nel78a]. computationally [TOO23]. Computer [Nel71b, Nel79b, NL95]. Computers [Nel75e, Nel80b]. Computing [NC71, Nel71b, Nel75a, Nel84, Nel93a, Nel80a]. Condition [Kel99]. Conditional [LN04]. Conferences [Nel83]. conjecture [Ham52, Uli58]. Conjugate [Nel94c]. Consensus [Paw04]. constants [IBV^{+}22]. Constrained [TFB21, AT18, Pri21]. constraints [Bre04, Lar97]. Construction [Nel68b, NCC^{+}70]. Contingency [Nel77b, Nel00c, Nel74a]. Continuous [DDM^{+}16, BC97, CS03, CS05, Nel98a]. contrast [Sel15]. Contrasts [Nel63b]. Contribution [Nel97]. control [Baş26]. controller [Baş26, LTJ95]. Controversies [Paw04]. Controversy [NLSS08]. Convergence [Gal22, Gal24, LRWW98, LRWW99, LPW12, McK98, McK99, SEABZ24, Tom95]. convergent [Cha12, PCB02]. Conversation [Sen03]. Cook [Nel98a]. Cooper [Nel90b]. coordination [LY12]. Correction [Ano61, ZZP^{+}22]. Correspondence [CNM00, Nel01a]. Corrigenda [Ait59]. Criticism [Cox04]. Cross [Nel79d]. Cross-Tabulated [Nel79d]. Crowder [Ano07]. crucial [Wes19]. cruise [Baş26]. Curve [Ano61, Nel61a]. curves [Nel92a].

D [Ano07, Nel64a, Nel98a]. Data [Nel68f, Nel69, Nel72c, Nel78a, Nel79c, NRCC91, Nel94a, NGMS94, LN00a, LN00b, Nel82, Nel92b, Nel94c, Pay04, Nel75b, Nel79d]. Dave [Ano07]. DDE [SEABZ24]. Decomposition [Nel85b, Mas07, WYWW15]. Decompositions [Gol69]. Deconstructing [NGL^{+}94]. Decrease [Kel99]. derived [Nel98c]. Description [NCC^{+}70]. Design [Nel56, Nel72b, Baş26]. designed [Bai04]. Designs [Nel62b, Nel68d]. Desmond [Car97, Les97, LR97, Nel97, Wir97]. Detection [Kel99, LN03b]. Developments [Nel84]. differentiable [DDM^{+}16]. differential [Mas06]. Digital [Nel80b]. Dimensional [Nel77b, Arn93, FBP19]. dimensionality [HN06]. Dimensions [LRWW98, LPW12, LRWW99, Lóc13, Meh20]. direct [Cha12, KiANS22]. direction [XD14]. directional [LY12]. discretized [WSKF13]. Discussion [Les97, LR97, Nel87b, NGL^{+}94, Nel97]. Dispersion [CCG12, NLB^{+}98]. dispersions [LN01]. distinctions [Van95]. distributed [KTL19]. distribution [RWJ13]. distributions [Arn93, Nel94c]. diversification [SEABZ24]. dose [Lar97]. Double [LN06a, LN00a]. double-exponential [LN00a]. Douglas [Nel01b]. Dr. [Nel90b].

each [Nel88]. earthquake [TA13]. ECG [WYWW15]. economic [SGK16].

Editor [BLN^{+}81, Nel98a, Nel01b, YYB^{+}76]. Editorial [Nel64b]. Editors [NS74, NKA92, NR97, Nel98b, NLSS08]. Edn [Che90]. Edward [Nel90b]. Effect [HN06, LN01, LN05, LN06b]. Effects [Boo07, LNP06, LNP17, Nel68a, Rob07, Dry10, NZE^{+}15, Gho08]. Efficient [Nel69, SS04]. Eigenvalue [TFB21]. elements [Mas09b]. enhanced [FZ06]. Enslein [Nel80b]. Enzyme [NRCC91]. Enzyme-Kinetic [NRCC91]. Epilogue [NC67]. Equation [Nel62a]. equations [Mas06, Mas09b, NL97]. Errata [NM65a]. error [Arn93]. Errors [NF84, BC97, Nel98a]. Essais [Nel64a]. Essence [Nel63a]. Estimating [NGN92, TA13, Des97a, Les97, LR97, Nel97, NL97, Wir97]. Estimation [Nel71c, RWJ13]. estimator [LN99]. estimators [LN03a]. Evaluation [Nel74b, Lar97]. examination [Nel04b]. Example [Nel79a]. expensive [TOO23]. Experimental [Nel56]. Experiments [Nel62b, Nel63b, NL95, Bai04, LN98, Nel65a, Nel65b, Nel72b]. expert [Nel88]. exponential [LN00a]. Extended [LN03a, NP87a, NL97, KTL19, LN00a, NP87b]. Extended-REML [LN03a]. extension [Bre13].

F [LR97, Nel94c, Nel97, Wir97]. Factor [Nel66, Nel77b, Nel00c]. Factorial [NL95]. Failure [NGMS94, Nel94c]. False [LN03b]. families [LN00a]. Fast [WYWW15]. Federer [Nel56]. Fiber [CCG12]. find [Lar97]. finite [Mas09b]. Fitting [Ano61, LN06b, Nel61a, Nel00a]. Flamant [Nel64a]. flight [Baş26]. flight-based [Baş26]. flow [NZE^{+}15]. Folks [Nel72c]. Foreword [Nel04a]. Form [Nel62a]. Formation [Nel69, Nel72a, NPT^{+}74, Nel75d]. Forms [Ait50, Ait59, Nel51]. Formulation [CCG12]. Fourier [WYWW15]. Fractional [Nel63b]. Frank [NL95]. Franklin [Nel77a]. FRS [ACHS04c, Ano07]. Function [Aki77, NP87a, ON75, FBP19, NM65a, NM65b, NP87b]. Functional [Nel00a]. Functions [Nel66, CS03, CS05, Des97a, Les97, LR97, Nel97, Wir97]. future [Nel93a]. Fuzzy [LL12, LTJ95].

G [Nel79d, Nel01b]. gaps [DCZ^{+}19]. Gaussian [HN55]. Geissler [LN00a]. General [CST04, Nel79c, Sen04, Nel65b]. Generalised [Sen04, LN01]. Generalization [Ano61, MJ10, Nel61a, Nel74a]. Generalized [LN98, LNP06, LNP17, MN83, MN89, Nel62a, NW72, Nel75a, NRCC91, LN96, LN06a, NGN92, Nel98c, Dry10, Nel86a, Pre84, Boo07, Bur91, Che90, Gho08, Lac90, Pla84, Rob07, Sch92, Wat86, Wil84]. generally [Nel68d]. Genetic [CS03, Mas09a, MM26, Mas06, Mas07, Mas09b, NZE^{+}15]. Gentle [Nel80a]. Geometric [MJ10]. George [Nel94c]. Gilding [NT02]. Given [CGVP15]. GLIM [Alt80, Col79, Nel75a, Nel85a]. GLMs [LN03b, Lee04]. global [CS03, CS05, SEABZ24]. globally [Cha12]. Gnanadesikan [Nel72b]. golden [NT02]. golden-section [NT02]. Graham [Nel79d]. Graphs [Nel76]. Graybill [Nel77a]. Grid [BPT06]. Group [Nel66]. growth [Nel61b]. Guide [Nel74b].

H [Nel80b]. Hammersley [Dvo57]. Hand [Ano07]. Hardware [YCH24]. Harris [ZZP^{+}22]. Hastings [Nel04b]. hawks [ZZP^{+}22]. heating [NZE^{+}15]. Heredity [Nel98d]. HGLMs [LNP11]. Hierarchical [LN96, LN01, Sen04, LN06a]. high [FBP19, Meh20]. high-dimensional [FBP19]. Honour [Ano07, ACHS04c]. Horton [Nel79c]. human [LN00a]. hybrid [CS05, LY12, RWJ13, SGK16, SEABZ24]. hybridized [CS03, LL12]. hyperbolic [Lóc13].

I/II [Lar97]. ICA [Sel15]. Identification [Nel63b]. II [Lar97, Nel65b]. IK [YCH24]. image [LL12]. implementation [SS04]. Implementing [GH12]. importance [Nel98b]. important [Nel93b]. Improved [Meh20, ZZP^{+}22]. Improvement [KiANS22, LN98, LNP11]. Improving [SEABZ24]. Incomplete [TFB21]. Independence [Ait50, Ait59, Nel51]. Index [CCG12]. inequality [AT18]. Inference [Paw04, Wak04, LN09a, LN09b]. information [Nel68d]. ingredients [Pay04]. initialization [TOO23, Wes19]. initio [PP23]. innovative [Pay04]. Input [NC71]. Input/Output [NC71]. integer [Bre13]. Interactive [Nel75a]. Interpretation [Nel85b, Nel54a]. Introduction [CCG12]. Invariance [Nel68e]. Inverse [Nel66, Nel68f, YCH24]. inversion [KiANS22]. isotropic [HN55]. iterations [SS01]. iterative [KiANS22].

J [Alt80, Ano07, Boo07, Bur91, Che90, Col79, Dal72, Lac90, Nel64a, Nel72b, Nel79d, Nel80a, Nel87a, Nel94c, Pla84, Pre84, Rob07, Sch92, Wat86, Wil84]. J. [Ait95, Bai04, Gow95, RTW95, Sea95]. Jacobian [YCH24]. JANM [YCH24]. JANM-IK [YCH24]. John [ACHS04c, Ano07, Dry10, Gar71, Gho08, Inn70, K.71, Nel63a, Nel78a, Van95, ACHS04b, Ano11, MGP11, Sen03, Sen04, Sen19]. Joint [NLB^{+}98, SGK16]. Joules [NZE^{+}15]. journals [Alt00, Nel01b]. Jr [Nel80a]. Julian [Ano11].

Kempthorne [Nel72c]. Kennedy [Nel80a]. kind [Nel77c]. Kinds [Nel62b]. Kinematics [YCH24]. Kinetic [NRCC91].

L [Nel75b, Nel79c, Nel98a]. Languages [Nel71b, Pay04]. Large [Rip04, eSAdSM18, Nel98c]. least [Nel74a]. Lee [Boo07, Dry10, Gho08, Rob07]. Lellouch [Nel64a]. Leroy [Nel72c]. Less [eSAdSM18]. Letter [Nel98a, Nel01b, NLSS08]. Letters [BLN^{+}81, NS74, NKA92, NR97, Nel98b, YYB^{+}76]. level [KBČ20]. Lévy [Baş26]. Likelihood [Lee04, LN05, LN09a, NP87a, NL92, Paw04, Des97a, LN00a, LNN07, Les97, LR97, Nel85a, Nel94c, Nel97, NL97, Nel00b, Wir97, LN09b, Rob07, Boo07, Dry10]. Likelihood-based [Lee04]. Likhachev [IBV^{+}22]. lily [NT02]. Limited [Aki77]. Linear [Boo07, Bur91, Che90, CST04, Gho08, Lac90, LNP06, LNP17, MN83, MN89, NW72, Nel75a, Nel77a, Nel79c, Nel82, NRCC91, Pla84, Pre84, Rob07, Sch92, Sen04, Wak04, Wat86, Wil84, Ait95, Bre04, Dry10, Gow95, LN96, LN98, LN01, LN06a, Nel74a, Nel77d, NGN92, Nel94d, Nel95a, Nel95b, Nel98c, RTW95, Sea95, Van95]. Lines [Nel90a]. localization [KNCS21]. Log [Nel74a]. logic [LTJ95]. Logistic [Ano61, Nel62a, Nel61a]. Longford [Nel94b]. longitudinal [BC97, Nel98a]. lot [SGK16]. Low [LRWW98, LRWW99, Lóc13].

M [Ano07, Nel78a, Nel87a, Nel94c]. MacLagan [Nel75c]. Main [Nel68a]. Makov [Nel94c]. Manual [Alt80]. Manuals [Nel78b]. marginal [LN04]. marginality [Nel98b, Nel00a]. Martin [Ano07]. material [PP23]. mathematical [Uli58]. Matrices [TFB21]. Matrix [Gol69]. Maxwell [Mas09b]. McCullagh [Bur91, Che90, Lac90, Pla84, Pre84, Sch92, Wil84, Wat86]. Mead} [PP23, Aki77, Arn93, AT18, Baş26, Bre04, Bre13, BPT06, Cha12, CS03, CS05, CCG12, CGVP15, DCZ^{+}19, DDM^{+}16, FBP19, FZ06, Gal22, Gal24, GH12, eSAdSM18, HNX03, HN06, IBV^{+}22, KTL19, Kel99, KiANS22, KBČ20, KNCS21, LRWW98, LRWW99, LPW12, Lar97, LTJ95, LL12, LY12, Lóc13, MM26, Mas06, Mas07, Mas09a, Mas09b, McK98, McK99, Meh20, MJ10, NZE^{+}15, NT02, ON75, OT09, Per24, Per26, PCB02, Pri21, PB26, RWJ13, SGK16, Sel15, SS01, SS04, SEABZ24, TOO23, TFB21, Tom95, TA13, WYWW15, Wes19, DW87, Wri12, WSKF13, XD14, YCH24, ZHZ^{+}07, ZZP^{+}22]. Mean [NLB^{+}98]. means [KTL19, LL12, NZE^{+}15]. measurement [ZHZ^{+}07]. mechanical [DCZ^{+}19]. medical [Alt00, Nel01b]. Medicine [Nel98a, Nel01b]. Memorian [NL95]. Mesh [AT18]. Mesh-based [AT18]. metamaterial [DCZ^{+}19]. Method [Aki77, CCG12, LRWW98, McK98, Bre13, Cha12, CS05, DCZ^{+}19, Gal22, eSAdSM18, HNX03, HN06, IBV^{+}22, KTL19, KiANS22, LRWW99, Lar97, Lóc13, Mas06, Mas07, McK99, NZE^{+}15, NM65a, NM65b, OT09, Pri21, RWJ13, Sel15, TOO23, WYWW15, Wri12, XD14, ZHZ^{+}07, ZZP^{+}22]. Methods [ACHS04c, Nel78a, Nel80b, Mas09a, Ano07]. Microcomputers [DW87]. Milton [Dal72, Gar71, Inn70, K.71]. Minimisation [Aki77]. Minimization [CCG12, ON75, TFB21, NM65a, NM65b, Sel15]. Missing [Nel54b]. Missing-Plot [Nel54b]. Mixed [CST04, NLSS08, Bre13]. mixed-integer [Bre13]. Mixed-Models [NLSS08]. Model [Cox04, CST04, Nel68e, Nel79c, IBV^{+}22, SNB^{+}01, Nel77a]. Model-Building [Nel68e]. Modeling [NLB^{+}98]. Modelling [Nel75a, Paw04, Wak04, Des97a, LN00a, LN00b, Nel97, Les97, LR97, Wir97]. Models [Ano07, Boo07, Gho08, Lee04, LNP06, LNP17, MN83, MN89, NW72, NF84, Nel86a, NRCC91, Nel92a, Nel98d, NLSS08, Rip04, Rob07, Sen04, ACHS04c, Ait95, BC97, Dry10, Gow95, LN96, LN98, LN01, LN04, LN05, LN06a, LN06b, LN09a, LN09b, Nel74a, Nel77d, Nel82, NGN92, Nel94d, Nel95a, Nel95b, Nel98c, Nel98a, RTW95, Sea95, Van95, Bur91, Che90, Lac90, Nel94b, Pla84, Pre84, Sch92, Wat86, Wil84]. modified [Arn93, DDM^{+}16, Lar97, Pri21]. Morgan [Nel87a]. most [Nel93b]. MR1229290 [Nel94c]. MR2757427 [LN09b]. MR2757428 [LN09b]. MR2757429 [LN09b]. MR2757430 [LN09b]. MR3363409 [NM65a].

Multi [Nel66, Nel68a, Nel77b, XD14]. Multi-Dimensional [Nel77b]. Multi-direction-based [XD14]. Multi-Factor [Nel66]. Multi-Way [Nel68a]. multidimensional [Mas07]. multiminima [CS03, CS05]. Multiresponse [Nel72b]. My [Nel77c].

N [Nel72b, Nel94b]. nanoparticle [ZHZ^{+}07]. Nearly [Nel90a]. negative [Nel54a]. neighbourhood [DDM^{+}16]. Nelder [ACHS04c, Ait95, Ano07, Ano11, Boo07, Dry10, Gho08, Gow95, PP23, Pre84, Rob07, RTW95, Sea95, Sen19, Wat86, Wri12, ACHS04b, Aki77, Alt80, Arn93, AT18, Bai04, Baş26, Bre04, Bre13, BPT06, Bur91, Cha12, Che90, CS03, CS05, CCG12, Col79, CGVP15, Dal72, DCZ^{+}19, DDM^{+}16, Dvo57, FBP19, FZ06, Gal22, Gal24, GH12, Gar71, eSAdSM18, Ham52, HNX03, HN06, Inn70, IBV^{+}22, K.71, KTL19, Kel99, KiANS22, KBČ20, KNCS21, Lac90, LRWW98, LRWW99, LPW12, Lar97, LTJ95, LL12, LY12, Lóc13, MM26, Mas06, Mas07, Mas09a, Mas09b, MGP11, McK98, McK99, Meh20, MJ10, NZE^{+}15, NT02, ON75, OT09, Per24, Per26, Pla84, PCB02, Pri21, PB26, RWJ13, SGK16, Sch92, Sel15, Sen03, Sen04, SS01, SS04, SEABZ24, TOO23, TFB21, Tom95, TA13]. Nelder [Uli58, Van95, WYWW15, Wes19, Wil84, DW87, WSKF13, XD14, YCH24, ZHZ^{+}07, ZZP^{+}22]. networks [KNCS21]. Newtonian [NZE^{+}15]. next [Nel93b]. Niall [Ano07]. no [Nel94c]. nodes [PP23, PP23]. Non [Wak04, DDM^{+}16, LN00b]. non-differentiable [DDM^{+}16]. Non-Linear [Wak04]. non-normal [LN00b]. Nonlinear [CCG12, Per24, Per26, Bre13, RWJ13]. nonorthogonal [Nel82]. Nonsmooth [Sel15]. Nonstationary [McK98, McK99]. Normal [Ait50, Ait59, LN00b]. North [Nel87a]. Note [CGVP15, Nel54b, Nel64b, Nel51, Nel61b, Nel62a]. null [Nel65a]. Numerical [KTL19].

Obituaries [Ano11]. Obituary [Nel90b]. October [Sen19]. Old [Gal24]. One [Nel77b, Nel00c]. Open [Gal24]. Optical [CCG12]. Optimal [Baş26, Des97a, Nel97, Lar97, PB26, Les97, LR97, Wir97]. Optimization [DW87, AT18, Bre04, Bre13, Cha12, CS03, CS05, DDM^{+}16, FBP19, FZ06, eSAdSM18, LY12, Mas07, Mas09a, Per24, Per26, Pri21, SEABZ24, TOO23, Tom95, ZZP^{+}22]. optimize [IBV^{+}22, SGK16]. optimizer [Baş26]. Optimizing [Arn93, LTJ95]. ordinary [Mas06]. organism [Nel61b]. Oriented [Nel79b]. origin [Nel92a]. Ornithology [Nel87a]. orthogonal [Nel65a, Nel65b]. Oscar [Nel72c]. other [Nel88, Wri12]. Output [NC71]. overcurrent [LY12]. overdispersion [LN00b].

P [Bur91, Che90, Lac90, Nel79b, Nel87a, Nel98a, Pla84, Sch92, Wat86, Wil84]. P. [Pre84]. Packages [Nel74b, Nel92b]. Pairwise [TFB21]. paper [Car97, Les97, LR97, Nel97, Wir97]. Papers [Bai04, Nel83]. Parabolic [CCG12]. Parabolic-Index [CCG12]. Parallel [Nel90a]. parallelisation [KBČ20]. Parameter [CGVP15, PB26]. parameters [GH12, Meh20, NGN92, TA13]. parsimony [LN03b]. particle [LY12]. Partitioned [MM26]. Party [Nel75a]. past [Nel93a]. patterns [Nel61b]. Pawitan [Dry10, Boo07, Gho08, Rob07]. Personal [Nel84]. Perspectives [CST04, Paw04]. perturbed [FBP19]. Phase [Lar97]. PIDD [Baş26]. Plackett [Nel75b]. Plot [Nel54b]. Plots [Nel90a]. Point [McK98, KTL19, McK99, NZE^{+}15]. points [Meh20]. Poisson [NF84]. Polynomials [Nel66, Nel68f, HNX03]. Position [Nel84]. Potential [Nel84]. Practical [TOO23]. Precision [Aki77]. Preface [ACHS04a]. Present [Nel84, Nel93a]. Presentation [Nel83]. Price [Nel79a]. Principle [Nel98d]. Principles [Bai04]. Probability [Nel72c]. problem [Dvo57, KTL19, MM26, Mas07, SGK16]. problem-solving [MM26]. problems [Bre13, LNN07, Mas09a, OT09, RWJ13, TOO23]. Procedure [ON75, FZ06, Per24, Per26]. process [HN55]. Professor [Ano07, ACHS04c]. Program [Nel75a]. Programming [NC67, NC71, RWJ13]. Proper [Wes19]. Properties [LRWW98, LRWW99]. pseudo [Nel00b]. pseudo-likelihood [Nel00b]. pseudolikelihood [NL92]. Pump [NGMS94, Nel94c]. Pump-Failure [NGMS94, Nel94c]. PY [PP23]. PY-Nodes [PP23]. Python [PP23].

Quadratic [Ait50, Ait59, Nel51]. quality [LN98, LNP11, Nel05]. quality-improvement [LN98]. Quantal [Nel68f]. Quantitative [Nel72b]. quantum [ZZP^{+}22]. Quasi [Nel85a, NP87a, Nel00b, Des97a, LN00a, Les97, LR97, NL92, Nel97, NL97, Wir97]. Quasi-Likelihood [NP87a, Nel85a, Nel00b, Des97a, LN00a, Les97, LR97, NL92, Nel97, NL97, Wir97]. quasilikelihood [LN99, NP87b]. Questions [Gal24, NGL^{+}94].

R [Alt80, Col79, Dal72, Nel64a, Nel72b, Nel75b, Nel79c]. Ralston [Nel80b]. Random [Boo07, Gho08, LNP06, LNP17, Rob07, Dry10, LN01, LN05, LN06b, Nel94b]. random-effect [LN01, LN05, LN06b]. randomized [Nel65a, Nel65b]. rapid [SEABZ24]. Rarities [Nel04b]. Rates [NF84]. ratio [LN00a]. Re [NGMS94, Nel94c]. Re-Analysis [NGMS94, Nel94c]. Reaction [NGMS94]. Reader [NGMS94]. real [Baş26, MM26]. real-coded [MM26]. Recherche [Nel72a, NPT^{+}74, Nel75d]. reformulation [Nel77d]. Regenerative [Aki77]. region [Lar97]. Regression [Nel68e, Nel68f, Nel85b, Wak04, BC97, Nel98a, Nel79a]. Rejoinder [Des97b, LN09b, Nel95a]. Relating [CGVP15]. relationship [LN00a]. relays [LY12]. Release [Alt80, Col79]. Remarkes [Cox04]. Remediation [Kel99]. REML [CST04, LN03a]. Replicates [Nel63b]. Research [Nel72a, NPT^{+}74, Nel75d, Nel93b]. Research/Formation [Nel72a, NPT^{+}74, Nel75d]. Residual [Nel90a]. Response [Nel66, Nel68f, Nel77b, NGMS94, Nel98d, Nel00c, NLSS08, FZ06, Nel00a]. Response-Surface [Nel98d, Nel00a]. restrained [BPT06]. Restricted [LPW12, Nel69]. Results [Gal24]. Review [Ano07, Boo07, Bur91, Col79, Dal72, Dry10, Gar71, Gho08, Inn70, K.71, Lac90, Nel56, Nel63a, Nel64a, Nel68c, Nel71a, Nel72b, Nel72c, Nel75b, Nel77a, Nel78a, Nel78b, Nel79c, Nel79a, Nel79b, Nel80b, Nel80a, Nel87a, Nel94b, Pla84, Pre84, Rob07, Sch92, Wat86, Wil84]. Reviews [Alt80, Che90, Nel79d]. RFIC [WSKF13]. Riemannian [Sel15]. Robert [Nel75c]. robustness [LN99]. roots [HNX03]. Round [Nel00c]. Rouquette [Nel64a]. Roy [Gar71, Inn70, K.71, Nel72b]. rules [Nel98b].

S [Nel72b, Nel79a, Nel79b, Nel80b]. same [Nel00b]. Sampling [HN55, SNB^{+}01]. Scaling [Nel76]. Scand [Nel94c]. scatter [SGK16]. scheme [KBČ20]. schemes [SNB^{+}01]. Schrödinger [Mas09b]. Schwartz [Nel64a]. Science [Nel86b, Nel99]. Sciences [Nel79c]. search [Cha12, CS05, DDM^{+}16, FZ06, NT02, Per24, Per26, SGK16, SS01, SS04, Wes19]. searching [PP23, WYWW15]. section [NT02]. see [Nel88]. seepage [OT09]. segmentation [LL12]. Selecting [Rip04, IBV^{+}22]. Selection [Nel98d]. Self [Aki77]. Semi [CCG12]. sensor [KNCS21]. sex [LN00a]. shark [Baş26]. should [Nel88]. signals [WYWW15]. Simple [Nel76, Nel61b]. Simplex [Aki77, CCG12, Gal24, LRWW98, McK98, ON75, DW87, Arn93, Cha12, CS05, FBP19, FZ06, GH12, HNX03, HN06, LRWW99, Lar97, LTJ95, LL12, Lóc13, McK99, NM65a, NM65b, OT09, PP23, Per24, Per26, PB26, RWJ13, Tom95, TA13, Wes19, Wri12, ZHZ^{+}07, ZZP^{+}22]. simplified [eSAdSM18]. simulated [TA13]. Simulation [ZHZ^{+}07, Cha12]. Singular [Nel85b, Mas07]. Singular-Value [Nel85b]. Sixty [Gal24]. size [ZHZ^{+}07]. sizing [SGK16]. smart [KNCS21]. Smith [Nel94c]. Social [Nel79c]. Solution [Mas09b, KTL19, Mas06, Mas07]. solutions [LNN07]. solve [SEABZ24]. solving [MM26, RWJ13]. Some [Cox04, Nel84, Alt00, Nel61b, NL92, Nel01b, Van95]. source [TA13]. Spacing [Nel62b, Meh20]. Splicing [Nel94a]. SPSS [Nel78b]. squares [Nel74a]. Srivastava [Nel72b]. Stagnation [Kel99, NZE^{+}15]. standardized [NGN92]. Stanley [Nel63a]. Static [KNCS21]. statics [KTL19]. Statist [Nel94c]. Statistical [Ait50, Ait59, Gol69, MN14, NC67, Nel69, NC71, Nel71b, Nel72a, NPT^{+}74, Nel74b, Nel75a, Nel75d, Nel76, Nel84, Nel92b, NGL^{+}94, Nel99, Paw04, Des97a, Les97, LR97, Nel51, Nel88, Nel93a, Nel93b, Nel97, Nel04b, Wir97, Dal72, Gar71, Inn70, K.71, Nel79b, Nel80b, Nel80a]. Statistics [Alt00, Ano07, Nel72c, Nel86b, Nel99, ACHS04c, Ait95, Gow95, Nel77c, Nel94d, Nel95a, Nel95b, RTW95, Sea95, Uli58, Van95, Nel98a, Nel01b, Nel71a, Nel87a]. Statistiques [Nel72a, NPT^{+}74, Nel75d]. steepest [Arn93]. Step [CCG12]. Stephens [Ano07]. Stochastic [Cha12, FZ06, Arn93, Tom95]. Storage [Nel69]. stretching [NZE^{+}15]. Strong [Nel98d]. structure [Nel65a, Nel65b]. structured [LN01]. structures [Pay04]. studies [BC97, Nel98a]. subject [Lar97]. subspace [KiANS22]. Sufficient [Kel99]. suitable [WSKF13]. Surface [Nel98d, FZ06, NZE^{+}15, Nel00a]. surfaces [Arn93]. SVD [Mas07]. swarm [LY12]. Synthesis [Nel05, LN01]. System [Alt80, Col79, Nel88]. Systematic [Nel62b]. Systems [Nel74b, KTL19].

T [Nel56, Nel87a, Nel94b]. Table [Nel68b, Nel68a, Nel77b]. Tables [Nel00c, Nel74a]. tabu [CS05]. Tabulated [Nel79d]. Technical [CGVP15]. Technology [Nel86b]. ten [Nel93b]. tenability [Van95]. tensors [Mas07]. Terms [Nel98d]. their [CGVP15]. theoretical [Nel61b]. Theory [Nel56, Nel77a]. thérapeutiques [Nel64a]. therapy [Lar97]. thing [Nel00b]. three [Arn93, KBČ20]. three-dimensional [Arn93]. three-level [KBČ20]. Topics [Nel71c]. toxicity [Lar97]. Training [Nel72a, NPT^{+}74, Nel75d]. Treatment [Nel65b]. trends [Alt00, Nel01b]. trial [Lar97]. Triangular [Nel69]. Two [LPW12, LN00b, Nel00c, KTL19]. two-point [KTL19].

U [Nel94c]. ultrafine [ZHZ^{+}07]. unbalanced [Nel92b]. unconfined [OT09]. Unconstrained [CCG12, eSAdSM18, Tom95]. underwater [KNCS21]. Unified [Boo07, Rob07, Dry10]. unobservables [LN09a, LN09b]. Unstable [Mas06]. Upton [Nel79d]. use [Lar97]. Useful [Nel66]. user [Nel88, Nel74b]. Using [Aki77, Kel99, KiANS22, KNCS21, LTJ95, PP23, Sel15].

Value [Nel85b, Mas07, Mas09a]. Values [CGVP15, Nel54b, PB26]. variable [DDM^{+}16]. Variance [Nel71c, Nel87b, Nel51, Nel54a, Nel65a, Nel65b]. variant [Lóc13, NT02, PCB02]. Variates [Ait50, Ait59]. variational [Mas09a]. vehicle [Baş26]. versus [TA13]. via [Baş26, Boo07, Dry10, FZ06, LN06b, Mas06, Mas07, Mas09b, Rob07]. View [Nel94a, LN04, LN09a, LN09b]. Views [Nel84]. visualization [SNB^{+}01]. Volkov [IBV^{+}22].

W [Nel80a]. Walter [Nel56]. wave [DCZ^{+}19]. Way [Nel68a]. ways [LN00b]. Weak [Nel98d]. Weak-Heredity [Nel98d]. Wedderburn [Nel75c]. Weighted [Nel68f]. white [Baş26]. Wilf [Nel80b]. William [Nel75c]. Working [Nel75a]. Works [ACHS04b].

Yates [NL95]. Years [Gal24, Nel93b]. Youngjo [Gho08, Dry10]. Yudi [Dry10, Gho08].

References

Adams:2004:P

[ACHS04a] N. M. Adams, M. J. Crowder, D. J. Hand, and D. A. Stephens. Preface. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Adams:2004:WJN

[ACHS04b] N. M. Adams, M. J. Crowder, D. J. Hand, and D. A. Stephens. The works of John Nelder. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Adams:2004:MMS

[ACHS04c] Niall Adams, M. J. (Martin J.) Crowder, D. J. (David J.) Hand, and David Stephens, editors. Methods and models in statistics: in honour of Professor John Nelder FRS. Imperial College Press, London, UK, 2004. ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). xiii + 246 pp. LCCN QA276.A1 M48 2004.

Aitken:1950:SIQ

[Ait50] A. C. Aitken. On the statistical independence of quadratic forms in normal variates. Biometrika, 37(1/2):93–96, June 1950. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332151. See note [Nel51] and

corrigenda [Ait59].

Aitken:1959:CSI

[Ait59] A. C. Aitken. Corrigenda: “On the Statistical Independence of Quadratic Forms in Normal Variates”. Biometrika, 46(1/2):279, June 1959. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332856. See

[Ait59].

Aitkin:1995:CJN

[Ait95] Murray Aitkin. Comments on J. A. Nelder: ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):85–86, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143934. See [Nel94d].

Akitt:1977:FMU

[Aki77] J. W. Akitt. Function minimisation using the Nelder and Mead simplex method with limited arithmetic precision: the self regenerative simplex. The Computer Journal, 20(1):84–85, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/20/1/84.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/84.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/85.tif.

Altman:1980:BRG

[Alt80] D. G. Altman. Book reviews: The GLIM System Release 3 Manual, by R. J. Baker and J. A. Nelder. Applied Statistics, 29(1): 98, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altman:2000:SMJ

[Alt00] Douglas G. Altman. Statistics in medical journals: some recent trends. Statistics in Medicine, 19(23):3275–3289, December 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [Nel01b].

Anonymous:1961:CFG

[Ano61] Anonymous. Correction: The fitting of a generalization of the logistic curve. Biometrics, 17(3):509, September 1961. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527846. See [Nel61a].

Anonymous:2007:BRBs

[Ano07] Anonymous. Book review: Methods and Models in Statistics, in Honour of Professor John Nelder, FRS by Niall M. Adams; Martin Crowder; D. J. Hand; Dave Stephens. Technometrics, 49(4): 501, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471414.

Anonymous:2011:OJA

[Ano11] Anonymous. Obituaries: John Ashworth Nelder; Julian Besag. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):499–504, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Arnholt:1993:OTD

[Arn93] Alan Todd Arnholt. Optimizing three-dimensional surfaces with stochastic error distributions: a comparison between the steepest ascent algorithm and the modified Nelder–Mead simplex algorithm. ProQuest LLC, Ann Arbor, MI, 1993. 160 pp.

URL http://gateway.proquest.com/openurl?url_ver=Z39.
88-2004&rft_val_fmt=info:ofi/fmt:kev:mtx:dissertation&
res_dat=xri:pqdiss&rft_dat=xri:pqdiss:9323103. Thesis

(Ph.D.)–University of Northern Colorado.

Audet:2018:MBN

[AT18] Charles Audet and Christophe Tribes. Mesh-based Nelder–Mead algorithm for inequality constrained optimization. Computational optimization and applications, 71(2):331–352, 2018. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Bailey:2004:PDE

[Bai04] R. A. Bailey. Principles of designed experiments in J. A. Nelder’s papers. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Basak:2026:ORP

[Baş26] Hasan Başak. Optimal real PIDD^{2} controller design for vehicle cruise control via Lévy flight-based white shark optimizer with Nelder–Mead algorithm. International Journal of Dynamics and Control, 14(1):Paper No. 36, 31, 2026.

Broemeling:1997:BAR

[BC97] Lyle D. Broemeling and Peyton Cook. A Bayesian analysis of regression models with continuous errors with application to longitudinal studies. Statistics in Medicine, 16(4):321–332, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [Nel98a].

Beale:1981:LE

[BLN^{+}81] E. M. L. Beale, P. W. Lane, J. A. Nelder, A. F. Bissell, and P. L. Goldsmith. Letters to the Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 30(3):231–234, September 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2988053.

Booth:2007:BRG

[Boo07] James G. Booth. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-likelihood, by Y. Lee, J. A. Nelder, and Y. Pawitan. Biometrics, 63(4):1296–1297, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burmen:2006:GRN

[BPT06] Árpád Bűrmen, Janez Puhan, and Tadej Tuma. Grid restrained Nelder–Mead algorithm. Computational optimization and applications, 34(3):359–375, 2006. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Brea:2004:NMO

[Bre04] Ebert Brea. Nelder–Mead optimization under linear constraints. PhD thesis, University of Southampton, Southampton, UK,

2004. ???? pp. URL https://www.proquest.com/pqdtglobal/

docview/305035514. Thesis (Ph.D.)–University of Southampton (United Kingdom).

Brea:2013:ENM

[Bre13] Ebert Brea. An extension of the Nelder–Mead method to nonlinear mixed-integer optimization problems. Revista Internacional de Métodos Numéricos para Cálculo y Diseño en Ingeniería, 29(3): 163–174, 2013.

Burridge:1991:BRB

[Bur91] Jim Burridge. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):361–364, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/2983054.

Carroll:1997:CPD

[Car97] R. J. Carroll. Comment on the paper by Desmond. Journal of Statistical Planning and Inference, 60(1):104–106, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0378375897900100. See [Des97a].

Choudhury:2012:SAF

[CCG12] Raja Roy Choudhury, Arundhati Roy Choudhury, and Mrinal Kanti Ghose. Semi analytical formulation of dispersion characteristics for a step and parabolic-index optical fiber: Introduction of Nelder–Mead simplex method for nonlinear unconstrained minimization. International Journal of Computer Applications, 49(??):38–44, July 2012. CODEN ???? ISSN 0975-

8887. URL https://www.ijcaonline.org/archives/volume49/
number22/7906-1292/.

Corominas:2015:TNR

[CGVP15] Albert Corominas, Alberto García-Villoria, and Rafael Pastor. Technical note: Relating to the parameter values given by Nelder and Mead in their algorithm. The Computer Journal, 58(1):157– 159, January 2015. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/58/1/157.

Chang:2012:SNM

[Cha12] Kuo-Hao Chang. Stochastic Nelder–Mead simplex method—a new globally convergent direct search method for simulation optimization. European Journal of Operational Research, 220(3):684–694, 2012. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Cheek:1990:BRG

[Che90] P. J. Cheek. Book reviews: Generalized Linear Models, 2nd Edn, by P. McCullagh and J. A. Nelder. Applied Statistics, 39(3):385– 386, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:2000:C

[CNM00] Geoff Clarke, J. A. Nelder, and Douglas C. Montgomery. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 49(1):107–110, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.jstor.

org/stable/2681260.

Collett:1979:BRG

[Col79] D. Collett. Book review: The GLIM System Release 3, by R. J. Baker, J. A. Nelder. Biometrics, 35(2):527–528, June 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2530359.

Cox:2004:SRM

[Cox04] D. R. Cox. Some remarkes on model criticism. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Chelouah:2003:GNM

[CS03] Rachid Chelouah and Patrick Siarry. Genetic and Nelder–Mead algorithms hybridized for a more accurate global optimization of continuous multiminima functions. European Journal of Operational Research, 148(2):335–348, 2003. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Chelouah:2005:HMC

[CS05] Rachid Chelouah and Patrick Siarry. A hybrid method combining continuous tabu search and Nelder–Mead simplex algorithms for the global optimization of multiminima functions. European Journal of Operational Research, 161(3):636–654, 2005. CODEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Cullis:2004:PAR

[CST04] B. R. Cullis, A. B. Smith, and R. Thompson. Perspectives of ANOVA, REML and a general linear mixed model. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Dalenius:1972:BRB

[Dal72] Tore Dalenius. Book review: Statistical Computation by R. C. Milton; J. A. Nelder. International Statistical Review = Revue Internationale de Statistique, 40(2):254–256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402775.

Dong:2019:AWB

[DCZ^{+}19] Jiaqi Dong, Weiqi Chen, Zhe Zeng, Qing-Hua Qin, and Yi Xiao. Analysis of wave band gaps in mechanical metamaterial based on Nelder–Mead method. Engineering Analysis with Boundary Elements, 103:109–115, 2019. ISSN 0955-7997 (print), 1873-197X (electronic).

Drazic:2016:CVN

[DDM^{+}16] Milan Dražić, Zorica Dražić, Nenad Mladenović, Dragan Urošević, and Qiu Hong Zhao. Continuous variable neighbourhood search with modified Nelder–Mead for non-differentiable optimization. IMA Journal of Management Mathematics, 27(1):75–88, 2016.

Desmond:1997:OEF

[Des97a] A. F. Desmond. Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):77–104, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378375896001231.

See discussion [Les97, Car97, Nel97, Wir97, LR97] and rejoinder [Des97b].

Desmond:1997:R

[Des97b] A. F. Desmond. Rejoinder. Journal of Statistical Planning and Inference, 60(1):116–121, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900161. See [Des97a].

Drygas:2010:BRY

[Dry10] Hilmar Drygas. Book review: Youngjo Lee, John A. Nelder, Yudi Pawitan: Generalized linear models with random effects: unified analysis via H-likelihood. Statistical Papers, 51(4):1007– 1008, December 2010. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00362-008-0175-7.

Dvoretzky:1957:PNH

[Dvo57] A. Dvoretzky. On a problem of Nelder and Hammersley. Bulletin of the Research Council of Israel. Section A. Mathematics, Physics and Chemistry, 6:115–118, 1957. ISSN 0366-2829.

Wouk:1987:OMN

[DW87] John E. Dennis, Jr. and Daniel J. Woods. Optimization on microcomputers: the Nelder–Mead simplex algorithm. In Arthur Wouk, editor, New Computing Environments: Microcomputers in Large-Scale Computing, pages 116–122. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1987.

Goncalves-e-Silva:2018:LMS

[eSAdSM18] Kayo Gonçalves e Silva, Daniel Aloise, Samuel Xavier de Souza, and Nenad Mladenović. Less is more: simplified Nelder–Mead method for large unconstrained optimization. Yugoslav Journal of Operations Research, 28(2):153–169, 2018. ISSN 0354-0243 (print), 1820-743X (electronic).

Fajfar:2019:NMS

[FBP19] Iztok Fajfar, Árpád Bűrmen, and Janez Puhan. The Nelder–Mead simplex algorithm with perturbed centroid for high-dimensional function optimization. Optimization Letters, 13(5):1011–1025, 2019. ISSN 1862-4472 (print), 1862-4480 (electronic).

Fan:2006:SRS

[FZ06] Shu-Kai S. Fan and Erwie Zahara. Stochastic response surface optimization via an enhanced Nelder–Mead simplex search procedure. Engineering Optimization, 38(1):15–36, 2006.

Galantai:2022:CNM

[Gal22] Aurél Galántai. Convergence of the Nelder–Mead method. Numerical Algorithms, 90(3):1043–1072, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01221-7.

Galantai:2024:NMS

[Gal24] Aurél Galántai. The Nelder–Mead simplex algorithm is sixty years old: New convergence results and open questions. Algorithms (Basel), 17(11):1–21, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-

4893/17/11/523.

Garside:1971:BRB

[Gar71] M. J. Garside. Book review: Statistical Computation, by Roy C. Milton; John A. Nelder. Journal of the Royal Statistical Society. Series A (General), 134(4):690–691, ???? 1971. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.jstor.org/stable/2343678.

Gao:2012:INM

[GH12] Fuchang Gao and Lixing Han. Implementing the Nelder–Mead simplex algorithm with adaptive parameters. Computational optimization and applications, 51(1):259–277, 2012. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Ghosh:2008:BRBb

[Gho08] Subir Ghosh. Book review: Generalized Linear Models with Random Effects by Youngjo Lee; John A. Nelder; Yudi Pawitan. Technometrics, 50(4):542–543, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471536.

Golub:1969:MDS

[Gol69] G. H. Golub. Matrix decompositions and statistical computation. In R. C. Milton and J. A. Nelder, editors, Statistical Computation, pages 365–397. Academic Press, New York, NY, USA, 1969.

Gower:1995:CJN

[Gow95] John Gower. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5 (2):91–92, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00143936. See [Nel94d].

Hammersley:1952:CN

[Ham52] J. M. Hammersley. On a conjecture of Nelder. Compositio Mathematica, 10:241–244, 1952. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

Hammersley:1955:SIG

[HN55] J. M. Hammersley and J. A. Nelder. Sampling from an isotropic Gaussian process. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51:652–662, 1955. CO-DEN PCPSA4. ISSN 0008-1981.

Han:2006:EDN

[HN06] Lixing Han and Michael Neumann. Effect of dimensionality on the Nelder–Mead simplex method. Optimization Methods and Software, 21(1):1–16, 2006. ISSN 1055-6788.

Han:2003:RCP

[HNX03] Lixing Han, Michael Neumann, and Jianhong Xu. On the roots of certain polynomials arising from the analysis of the Nelder–Mead simplex method. Linear Algebra and its Applications, 363(C):109– 124, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print),

1873-1856 (electronic). URL http://www.elsevier.nl/gej-ng/
10/30/19/212/25/35/abstract.html. Special issue on nonnega-

tive matrices, M-matrices and their generalizations (Oberwolfach, 2000).

Ivanov:2022:ANM

[IBV^{+}22] A. M. Ivanov, F. S. Belyaev, A. E. Volkov, S. P. Belyaev, and N. N. Resnina. Application of the Nelder–Mead method to optimize the selecting of the Likhachev–Volkov model constants. Vestnik Sankt-Peterburgskogo Universiteta. Matematika. Mekhanika. Astronomiya, 9(67)(4):693–704, 2022.

Innis:1970:BRS

[Inn70] George Innis. Book review: Statistical Computation, by Roy C. Milton, John A. Nelder. Biometrics, 26(4):871, December 1970.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528747.

K:1971:BRBe

[K.71] G. G. K. Book review: Statistical Computation by Roy C. Milton; John A. Nelder. Technometrics, 13(3):710–711, August 1971. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1267191.

Kriauziene:2020:TLP

[KBČ20] Rima Kriauzienė, Andrej Bugajev, and Raimondas Čiegis. A three-level parallelisation scheme and application to the Nelder– Mead algorithm. Mathematical Modelling and Analysis, 25(4):584– 607, 2020.

Kelley:1999:DRS2

[Kel99] C. T. Kelley. Detection and remediation of stagnation in the Nelder–Mead algorithm using a sufficient decrease condition. SIAM Journal on Optimization, 10(1):43–55, October/November 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (elec-

tronic). URL http://epubs.siam.org/sam-bin/dbq/article/
31520.

Kitaoka:2022:INM

[KiANS22] Haru Kitaoka, Ken ichi Amano, Naoya Nishi, and Tetsuo Sakka. Improvement of the Nelder–Mead method using direct inversion in iterative subspace. Optimization and Engineering, 23(2):1033– 1055, 2022. ISSN 1389-4420 (print), 1573-2924 (electronic).

Kulandaivel:2021:SLU

[KNCS21] Madhumitha Kulandaivel, Arulanand Natarajan, Bharathi Priya Chandrasekaran, and Anandamurugan Selvaraj. Static localization for underwater acoustics sensor networks using Nelder–Mead algorithm for smart cities. Computational Intelligence, 37(4):1691– 1705, November 2021. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

Kalyukh:2019:NST

[KTL19] Yu. I. Kalyukh, A. N. Trofimchuk, and A. G. Lebed’. Numerical solution of a two-point statics problem for distributed extended systems by means of the Nelder–Mead method. Natsional’ naya Akademiya Nauk Ukrainy. Institut Kibernetiki im. V. M. Glushkova. Kibernetika i Sistemnyĭ Analiz, 55(4):109–118, 2019.

Lachenbruch:1990:BRG

[Lac90] P. A. Lachenbruch. Book review: Generalized Linear Models, by P. McCullagh, J. A. Nelder. Biometrics, 46(4):1231–1232, December 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2532465.

Larsen:1997:EUN

[Lar97] Lois M. Larsen. An evaluation of the use of the Nelder–Mead modified simplex method to find the region of optimal dose therapy combinations subject to toxicity constraints in a Phase I/II clinical trial. Thesis (Ph.D.), University of Northern Colorado, Greeley, CO, USA, 1997. 196 pp. URL https://www.proquest.

com/pqdtglobal/docview/304369251.

Lee:2004:LBM

[Lee04] Y. Lee. Likelihood-based models beyond GLMs. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Lesperance:1997:DDP

[Les97] Mary L. Lesperance. Discussion on A. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):106– 108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900112. See [Des97a].

Li:2012:FMI

[LL12] Yan Ling Li and Ting Liu. Fuzzy means image segmentation based on hybridized Nelder–Mead simplex algorithm. Mathematics in Practice and Theory. Shuxue de Shijian yu Renshi, 42(14):218– 223, 2012. ISSN 1000-0984.

Lee:1996:HGL

[LN96] Y. Lee and J. A. Nelder. Hierarchical generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):619–678, 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://links.jstor.org/sici?sici=0035-9246(
1996)58:4<619:HGLM>2.0.CO;2-8&origin=MSN. With discus-

sion.

Lee:1998:GLM

[LN98] Y. Lee and J. A. Nelder. Generalized linear models for the analysis of quality-improvement experiments. The Canadian Journal of Statistics = La revue canadienne de statistique, 26(1):95–105, 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1999:RQE

[LN99] Y. Lee and John A. Nelder. The robustness of the quasilikelihood estimator. The Canadian Journal of Statistics = La revue canadienne de statistique, 27(2):321–327, 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:2000:RBD

[LN00a] Y. Lee and J. A. Nelder. The relationship between doubleexponential families and extended quasi-likelihood families, with application to modelling Geissler’s human sex ratio data. Applied Statistics, 49(3):413–419, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2000:TWM

[LN00b] Y. Lee and J. A. Nelder. Two ways of modelling overdispersion in non-normal data. Applied Statistics, 49(4):591–598, 2000. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2001:HGL

[LN01] Youngjo Lee and John A. Nelder. Hierarchical generalised linear models: a synthesis of generalised linear models, random-effect models and structured dispersions. Biometrika, 88(4):987–1006, 2001. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Lee:2003:ERE

[LN03a] Youngjo Lee and John A. Nelder. Extended-REML estimators. Journal of Applied Statistics, 30(8):845–856, 2003. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2003:FPD

[LN03b] Youngjo Lee and John A. Nelder. False parsimony and its detection with GLMs. Journal of Applied Statistics, 30(5):477–483, 2003. ISSN 0266-4763 (print), 1360-0532 (electronic).

Lee:2004:CMM

[LN04] Youngjo Lee and John A. Nelder. Conditional and marginal models: another view. Statistical Science, 19(2):219–238, 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2005:LRE

[LN05] Youngjo Lee and John A. Nelder. Likelihood for random-effect models. SORT, 29(2):141–182, 2005. With a discussion and a rejoinder by the authors.

Lee:2006:DHG

[LN06a] Youngjo Lee and John A. Nelder. Double hierarchical generalized linear models. Applied Statistics, 55(2):139–185, 2006. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:2006:FAR

[LN06b] Youngjo Lee and John A. Nelder. Fitting via alternative randomeffect models. Statistics and Computing, 16(1):69–75, 2006. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lee:2009:LIM

[LN09a] Youngjo Lee and John A. Nelder. Likelihood inference for models with unobservables: another view. Statistical Science, 24(3):255– 269, 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2009:RLI

[LN09b] Youngjo Lee and John A. Nelder. Rejoinder: Likelihood inference for models with unobservables another view [MR2757427; MR2757428; MR2757430; MR2757429]. Statistical Science, 24(3): 294–302, 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Lee:2007:LPS

[LNN07] Youngjo Lee, John A. Nelder, and Maengseok Noh. H-likelihood: problems and solutions. Statistics and Computing, 17(1):49–55, 2007. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Lee:2006:GLM

[LNP06] Youngjo Lee, John A. Nelder, and Yudi Pawitan. Generalized Linear Models with Random Effects, volume 106 of Monographs on Statistics and Applied Probability. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-631-5; 1-58488-631-5. x + 396 pp. Unified analysis via H-likelihood, With 1 CD-ROM (Windows).

Lee:2011:HQI

[LNP11] Youngjo Lee, John A. Nelder, and Heejin Park. HGLMs for quality improvement. Applied Stochastic Models in Business and Industry, 27(3):315–328, 2011.

Lee:2017:GLM

[LNP17] Youngjo Lee, John A. Nelder, and Yudi Pawitan. Generalized Linear Models with Random Effects, volume 153 of Monographs on Statistics and Applied Probability. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2017. ISBN 1-4987-2061-7. xix + 446 pp. Unified analysis via H-likelihood.

Locsi:2013:HVN

[Lóc13] Levente Lócsi. A hyperbolic variant of the Nelder–Mead simplex method in low dimensions. Acta Universitatis Sapientiae. Mathematica. An International Scientific Journal of Sapientia University, 5(2):169–183, 2013.

Lagarias:2012:CRN

[LPW12] Jeffrey C. Lagarias, Bjorn Poonen, and Margaret H. Wright. Convergence of the restricted Nelder–Mead algorithm in two dimensions. SIAM Journal on Optimization, 22(2):501–532, ???? 2012. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Liang:1997:DFD

[LR97] Kung-Yee Liang and Paul J. Rathouz. Discussion on A. F. Desmond’s paper: Optimal estimating functions, quasilikelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):113–116, April 30, 1997. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375897900148. See [Des97a].

Lagarias:1998:CPN

[LRWW98] Jeffrey C. Lagarias, James A. Reeds, Margaret H. Wright, and Paul E. Wright. Convergence properties of the Nelder–Mead simplex method in low dimensions. SIAM Journal on Optimization, 9(1):112–147, October/December 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/30347.

Lagarias:1999:CPN

[LRWW99] Jeffrey C. Lagarias, James A. Reeds, Margaret H. Wright, and Paul E. Wright. Convergence properties of the Nelder–Mead simplex method in low dimensions. SIAM Journal on Optimization, 9(1):112–147, 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Li:1995:OFL

[LTJ95] Wei Li, Qunhua Tan, and Hartmut Janocha. Optimizing fuzzy logic controller using Nelder and Mead’s simplex algorithm. Control Theory & Applications. Kongzhi Lilun yu Yingyong, 12(5): 616–622, 1995. ISSN 1000-8152.

Liu:2012:NHN

[LY12] An Liu and Ming-Ta Yang. A new hybrid Nelder–Mead particle swarm optimization for coordination optimization of directional overcurrent relays. Mathematical Problems in Engineering, pages 1–18, 2012.

Mastorakis:2006:UOD

[Mas06] Nikos E. Mastorakis. Unstable ordinary differential equations: solution via genetic algorithms and the method of Nelder–Mead. WSEAS Transactions on Mathematics, 5(12):1276–1281, 2006. ISSN 1109-2769 (print), 2224-2880 (electronic).

Mastorakis:2007:SVD

[Mas07] Nikos E. Mastorakis. The singular value decomposition (SVD) in tensors (multidimensional arrays) as an optimization problem. Solution via genetic algorithms and method of Nelder–Mead. WSEAS Transactions on Systems, 6(1):17–23, 2007. ISSN 1109-2777.

Mastorakis:2009:GAN

[Mas09a] Nikos E. Mastorakis. Genetic algorithms with Nelder–Mead optimization in the variational methods of boundary value problems. WSEAS Transactions on Mathematics, 8(3):107–116, 2009. ISSN 1109-2769 (print), 2224-2880 (electronic).

Mastorakis:2009:SSM

[Mas09b] Nikos E. Mastorakis. Solution of the Schrödinger–Maxwell equations via finite elements and genetic algorithms with Nelder–Mead. WSEAS Transactions on Mathematics, 8(4):169–176, 2009. ISSN 1109-2769 (print), 2224-2880 (electronic).

McKinnon:1998:CNM

[McK98] K. I. M. McKinnon. Convergence of the Nelder–Mead simplex method to a nonstationary point. SIAM Journal on Optimization, 9(1):148–158, October/December 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/30348.

McKinnon:1999:CNM

[McK99] K. I. M. McKinnon. Convergence of the Nelder–Mead simplex method to a nonstationary point. SIAM Journal on Optimization, 9(1):148–158, 1999. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Mehta:2020:INM

[Meh20] V. K. Mehta. Improved Nelder–Mead algorithm in high dimensions with adaptive parameters based on Chebyshev spacing points. Engineering Optimization, 52(10):1814–1828, 2020.

McCullagh:2011:JAN

[MGP11] Peter McCullagh, Bob Gilchrist, and Roger Payne. John Ashworth Nelder, 1924–2010. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174(2):499–502, April 2011. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL

http://www.jstor.org/stable/23014410.

Moraglio:2010:GGN

[MJ10] Alberto Moraglio and Colin G. Johnson. Geometric generalization of the Nelder–Mead algorithm. In Evolutionary computation in combinatorial optimization, volume 6022 of Lecture Notes in Comput. Sci., pages 190–201. Springer, Berlin, 2010. ISBN 3-642-12138-1; 3-642-12138-1.

Majhi:2026:PGA

[MM26] Neha Majhi and Rajashree Mishra. Partitioned genetic algorithm with Nelder–Mead for real-coded problem-solving and it’s application. In Proceedings of the International Conference on Computational Mathematics and Applications, volume 1520 of Lect. Notes Netw. Syst., pages 169–182. Springer, Singapore, 2026. ISBN 981-9689-86-4; 981-9689-87-2.

McCullagh:1983:GLM

[MN83] P. McCullagh and J. A. Nelder. Generalized Linear Models. Monographs on Statistics and Applied Probability. Chapman and Hall, Ltd., London, UK, 1983. ISBN 0-412-23850-0. xiii + 261 pp. LCCN H61.25; QA276.M38.

McCullagh:1989:GLM

[MN89] P. McCullagh and J. A. Nelder. Generalized Linear Models. Monographs on Statistics and Applied Probability. Chapman and Hall, Ltd., London, UK, second edition, 1989. ISBN 0-412-31760-5. xix + 511 pp.

Milton:2014:SC

[MN14] Roy C. Milton and John A. Nelder, editors. Statistical Computation. Elsevier Science and Technology, 2014. ISBN 1-4832-5802-5. LCCN QA276.4 S738 1969.

Nelder:1967:SPE

[NC67] J. A. Nelder and B. E. Cooper. Statistical programming: Epilogue. Applied Statistics, 16(2):149–151, 1967. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:SCI

[NC71] J. A. Nelder and B. E. Cooper. Statistical computing 1970: Input/ output in statistical programming. Applied Statistics, 20(1):56– 73, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1970:CDA

[NCC^{+}70] J. A. Nelder, B. E. Cooper, J. M. Craddock, J. C. Gower, P. J. Harrison, and I. D. Hill. The construction and description of algorithms. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):255–260, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.jstor.

org/stable/2986837.

Nelder:1951:NSI

[Nel51] J. A. Nelder. A note on the statistical independence of quadratic forms in the analysis of variance. Biometrika, 38:482–483, 1951.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nelder:1954:INC

[Nel54a] J. A. Nelder. The interpretation of negative components of variance. Biometrika, 41:544–548, 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nelder:1954:NMP

[Nel54b] J. A. Nelder. A note on missing-plot values. Biometrics, 10(3): 400–401, September 1954. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3001594.

Nelder:1956:BRB

[Nel56] J. A. Nelder. Book review: Experimental Design, Theory and Application, by Walter T. Federer. Journal of the Royal Statistical Society. Series A (General), 119(3):340–341, ???? 1956. CO-DEN JSSAEF. ISSN 0035-9238. URL https://www.jstor.org/

stable/2342744.

Nelder:1961:FGL

[Nel61a] J. A. Nelder. The fitting of a generalization of the logistic curve. Biometrics, 17:89–110, 1961. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Nelder:1961:NSG

[Nel61b] J. A. Nelder. A note on some growth patterns in a simple theoretical organism. Biometrics, 17:220–228, 1961. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Nelder:1962:NAF

[Nel62a] J. A. Nelder. 182. Note: an alternative form of a generalized logistic equation. Biometrics, 18(4):614–616, December 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527907.

Nelder:1962:NKS

[Nel62b] J. A. Nelder. New kinds of systematic designs for spacing experiments. Biometrics, 18(3):283–307, September 1962. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2527473.

Nelder:1963:BRB

[Nel63a] J. A. Nelder. Book review: The Essence of Biometry, by John Stanley. Journal of the Royal Statistical Society. Series A (General), 126(4):600–601, ???? 1963. CODEN JSSAEF. ISSN 0035-

9238. URL https://www.jstor.org/stable/2982590.

Nelder:1963:ICF

[Nel63b] J. A. Nelder. Identification of contrasts in fractional replicates of 2^{n} experiments. Applied Statistics, 12(1):38–43, March 1963. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1964:BRE

[Nel64a] J. A. Nelder. Book review: Les Essais thérapeutiques cliniques, by D. Schwartz, R. Flamant, J. Lellouch, C. Rouquette. Biometrics, 20(2):376, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/

2528410. Special issue: In memoriam: Ronald Aylmer Fisher, 1890–1962.

Nelder:1964:EN

[Nel64b] J. A. Nelder. Editorial note. Biometrics, 20(4):883, December 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528138.

Nelder:1965:AREa

[Nel65a] J. A. Nelder. The analysis of randomized experiments with orthogonal block structure. I. Block structure and the null analysis of variance. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 283:147–162, 1965. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Nelder:1965:AREb

[Nel65b] J. A. Nelder. The analysis of randomized experiments with orthogonal block structure. II. Treatment structure and the general analysis of variance. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 283:163–178, 1965. CO-DEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Nelder:1966:IPU

[Nel66] J. A. Nelder. Inverse polynomials, a useful group of multi-factor response functions. Biometrics, 22(1):128–141, March 1966. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528220.

Nelder:1968:AAM

[Nel68a] J. A. Nelder. Algorithms: Algorithm AS 8: Main effects from a multi-way table. Applied Statistics, 17(3):277–279, September 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.edu/apstat/8.

Nelder:1968:AAC

[Nel68b] J. A. Nelder. Algorithms: Algorithm AS 9: Construction of additive table. Applied Statistics, 17(3):279–283, September 1968. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/apstat/9.

Nelder:1968:BR

[Nel68c] J. A. Nelder. Book review. The Computer Journal, 10(4):367, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/367.full.pdf+html.

Nelder:1968:CIG

[Nel68d] J. A. Nelder. The combination of information in generally balanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 30:303–311, 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://links.jstor.org/sici?sici=0035-9246(
1968)30:2<303:TCOIIG>2.0.CO;2-D&origin=MSN.

Nelder:1968:RMB

[Nel68e] J. A. Nelder. Regression, model-building and invariance. Journal of the Royal Statistical Society. Series A (General), 131(3):303– 329, ???? 1968. CODEN JSSAEF. ISSN 0035-9238. URL https:

//www.jstor.org/stable/2343525.

Nelder:1968:WRQ

[Nel68f] J. A. Nelder. Weighted regression, quantal response data, and inverse polynomials. Biometrics, 24(4):979–985, December 1968. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528884.

Nelder:1969:SAA

[Nel69] J. A. Nelder. Statistical algorithms: Algorithm AS 20: The efficient formation of a triangular array with restricted storage for data. Applied Statistics, 18(2):203–206, June 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/20.

Nelder:1971:BRB

[Nel71a] J. A. Nelder. Book review: Statistics In Biology, by Chester I. Bliss. Journal of the Royal Statistical Society. Series A (General), 134(2):246–247, ???? 1971. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2343879.

Nelder:1971:SCS

[Nel71b] J. A. Nelder. Statistical computing 1970: Statistical computing and computer languages. Applied Statistics, 20(1):25–32, 1971. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1971:TVC

[Nel71c] J. A. Nelder. [Topics in variance component estimation]. Biometrics, 27(3):750, September 1971. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.

jstor.org/stable/2528616.

Nelder:1972:STR

[Nel72a] J. Nelder. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 40(3):384–388, December 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402476.

Nelder:1972:BRBa

[Nel72b] J. A. Nelder. Book review: Analysis and Design of Certain Quantitative Multiresponse Experiments by S. N. Roy; R. Gnanadesikan; J. N. Srivastava. International Statistical Review = Revue Internationale de Statistique, 40(2):256, August 1972. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402776.

Nelder:1972:BRBb

[Nel72c] J. A. Nelder. Book review: Probability, Statistics and Data Analysis, by Oscar Kempthorne; Leroy Folks. Journal of the Royal Statistical Society. Series A (General), 135(3):430–431, ???? 1972. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2344621.

Nelder:1974:LLM

[Nel74a] J. A. Nelder. Log linear models for contingency tables: a generalization of classical least squares. Applied Statistics, 23:323–329, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1974:UGE

[Nel74b] J. A. Nelder. A User’s guide to the evaluation of statistical packages and systems. International Statistical Review = Revue Internationale de Statistique, 42(3):291–298, December 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402986.

Nelder:1975:AWP

[Nel75a] J. A. Nelder. Announcement by the working party on statistical computing: GLIM (generalized linear interactive modelling program). Applied Statistics, 24(2):259–261, 1975. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1975:BRB

[Nel75b] J. A. Nelder. Book review: The Analysis of Categorical Data, by R. L. Plackett. Journal of the Royal Statistical Society. Series A (General), 138(1):99–100, ???? 1975. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2345253.

Nelder:1975:RWM

[Nel75c] J. A. Nelder. Robert William MacLagan Wedderburn, 1947–1975. Journal of the Royal Statistical Society. Series A (General), 138 (4):587, ???? 1975. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2345239.

Nelder:1975:STR

[Nel75d] J. A. Nelder. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 43(2):227–232, August 1975. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1402904.

Nelder:1975:CB

[Nel75e] John A. Nelder. Computers in Biology. The Wykeham science series. Wykeham Publications, London, UK, 1975. ISBN 0-387-91120-0. viii + 158 pp. LCCN QH324.2 .N44 1975.

Nelder:1976:SAA

[Nel76] J. A. Nelder. Statistical algorithms: Algorithm AS 96: a simple algorithm for scaling graphs. Applied Statistics, 25(1):94–96, March 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-

9876 (electronic). URL http://lib.stat.cmu.edu/apstat/96.

Nelder:1977:BRB

[Nel77a] J. A. Nelder. Book review: Theory and Application of the Linear Model, by Franklin A. Graybill. Journal of the Royal Statistical Society. Series A (General), 140(3):384–385, ???? 1977. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344939.

Nelder:1977:MDC

[Nel77b] J. A. Nelder. Multi-dimensional contingency table with one factor as a response. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):41–42, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2988218.

Nelder:1977:MKS

[Nel77c] J. A. Nelder. My kind of statistics. Bulletin in Applied Statistics (BIAS), 4(2):77–81, 1977. CODEN ???? ISSN ????

Nelder:1977:RLM

[Nel77d] J. A. Nelder. A reformulation of linear models. Journal of the Royal Statistical Society. Series A (General), 140(1):48–76, 1977. CODEN JSSAEF. ISSN 0035-9238. With discussion.

Nelder:1978:BRB

[Nel78a] J. A. Nelder. Book review: Computational Methods for Data Analysis, by John M. Chambers. Journal of the Royal Statistical Society. Series A (General), 141(4):550–551, ???? 1978. CODEN JS-

SAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2344494.

Nelder:1978:RMB

[Nel78b] J. A. Nelder. A review of the manuals for BMDP and SPSS: Comment. Journal of the American Statistical Association, 73 (361):89–90, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2286523.

Nelder:1979:BRR

[Nel79a] J. A. Nelder. Book review: Regression Analysis by Example, by S. Chatterjee, B. Price. Biometrics, 35(1):355–356, March 1979. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2529957.

Nelder:1979:BRS

[Nel79b] J. A. Nelder. Book review: Statistical Analysis, a Computer Oriented Approach, by A. A. Afifi, S. P. Azen. Biometrics, 35(3):703, September 1979. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
2530266.

Nelder:1979:BRB

[Nel79c] J. A. Nelder. Book review: The General Linear Model, Data Analysis in the Social and Behavioural Sciences, by R. L. Horton. Journal of the Royal Statistical Society. Series A (General), 142(3):385, ???? 1979. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2982495.

Nelder:1979:BRA

[Nel79d] J. A. Nelder. Book reviews: The Analysis of Cross-Tabulated Data, by Graham J. G. Upton. Applied Statistics, 28(3):306–307, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1980:BRS

[Nel80a] J. A. Nelder. Book review: Statistical Computing, by W. J. Kennedy, Jr., J. E. Gentle. Biometrics, 36(4):743–744, December 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/2556137.

Nelder:1980:BRB

[Nel80b] J. A. Nelder. Book review: Statistical Methods for Digital Computers by K. Enslein; A. Ralston; H. S. Wilf. International Statistical Review = Revue Internationale de Statistique, 48(1):120–121, April 1980. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823

(electronic). URL http://www.jstor.org/stable/1402419.

Nelder:1982:LMN

[Nel82] J. A. Nelder. Linear models and nonorthogonal data. Utilitas Mathematica, 21:141–152, 1982. CODEN UTMADA. ISSN 0315-3681. Special issue dedicated to Frank Yates on the occasion of his eightieth birthday, Vol. B.

Nelder:1983:PPC

[Nel83] John A. Nelder. Presentation of papers at conferences. Biometrics, 39(3):799, September 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2531116.

Nelder:1984:PPP

[Nel84] J. A. Nelder. Present position and potential developments: Some personal views: Statistical computing. Journal of the Royal Statistical Society. Series A (General), 147(2):151–160, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2981672.

Nelder:1985:QLG

[Nel85a] J. A. Nelder. Quasi-likelihood and GLIM. In Generalized linear models (Lancaster, 1985), volume 32 of Lect. Notes Stat., pages 120–127. Springer, Berlin, 1985. ISBN 3-540-96224-7.

Nelder:1985:AIS

[Nel85b] John A. Nelder. An alternative interpretation of the singularvalue decomposition in regression. The American Statistician, 39 (1):63–64, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683911.

Nelder:1986:GAM

[Nel86a] J. A. Nelder. [Generalized additive models]: Comment. Statistical Science, 1(3):312, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013606.

Nelder:1986:SST

[Nel86b] J. A. Nelder. Statistics, science and technology. Journal of the Royal Statistical Society. Series A (General), 149(2):109–121, ???? 1986. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981525.

Nelder:1987:BRS

[Nel87a] J. A. Nelder. Book review: Statistics in Ornithology, by B. J. T. Morgan, P. M. North. Biometrics, 43(1):253–254, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2531971.

Nelder:1987:DWA

[Nel87b] J. A. Nelder. Discussion: What is an analysis of variance? Annals of Statistics, 15(3):930–931, September 1987. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176350480.

Nelder:1988:HSS

[Nel88] J. A. Nelder. How should the statistical expert system and its user see each other? In COMPSTAT 1988 (Copenhagen, 1988), pages 107–116. Physica, Heidelberg, West Germany, 1988. ISBN 3-7908-0411-8.

Nelder:1990:NPL

[Nel90a] J. A. Nelder. Nearly parallel lines in residual plots. The American Statistician, 44(3):221–222, August 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2685341.

Nelder:1990:ODB

[Nel90b] J. A. Nelder. Obituary: Dr. Brian Edward Cooper, 1935–90. Journal of the Royal Statistical Society. Series A (General), 153(3):397, 1990. CODEN JSSAEF. ISSN 0035-9238.

Nelder:1992:MCT

[Nel92a] J. A. Nelder. Models for curves through the origin. Journal of Applied Statistics, 19(4):551–552, 1992. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:1992:SPU

[Nel92b] John A. Nelder. Statistical packages and unbalanced data. Computational Statistics & Data Analysis, 14(3):403–406, October 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352

(electronic). URL http://www.sciencedirect.com/science/
article/pii/016794739290050P.

Nelder:1993:ASC

[Nel93a] J. A. Nelder. Aspects of statistical computing, past, present and future. Statistica Neerlandica, 47(1):3–8, March 1993. CO-DEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic).

URL https://onlinelibrary.wiley.com/doi/epdf/10.1111/
j.1467-9574.1993.tb01402.x.

Nelder:1993:MIA

[Nel93b] J. A. Nelder. The most important areas of statistical research in the next ten years. Statistics and Computing, 3(4): 202–203, December 1993. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00141783.

Nelder:1994:AVS

[Nel94a] J. A. Nelder. An alternative view of the splicing data. Applied Statistics, 43(3):469–476, 1994. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1994:BRB

[Nel94b] J. A. Nelder. Book review: Random Coefficient Models, by N. T. Longford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):507, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983541.

Nelder:1994:RAP

[Nel94c] J. A. Nelder. A re-analysis of the pump-failure data. Comment on: “Conjugate likelihood distributions” [Scand. J. Statist. 20 (1993), no. 2, 147–156; MR1229290 (94g:62007)] by E. I. George, U. E. Makov and A. F. M. Smith. Scandinavian Journal of Statistics. Theory and Applications, 21(2):187–191, 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). With a reply by George, Makov and Smith.

Nelder:1994:SLM

[Nel94d] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 4(4):221–234, December 1994. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00156745. See comments [Nel95b, Ait95, Gow95, Van95, RTW95, Sea95] and rejoinder [Nel95a].

Nelder:1995:RCS

[Nel95a] J. A. Nelder. Rejoinder to comments on ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):109–111, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-

1375 (electronic). URL http://link.springer.com/article/
10.1007/BF00143940. See [Nel94d].

Nelder:1995:SLM

[Nel95b] J. A. Nelder. The statistics of linear models: back to basics. Statistics and Computing, 5(2):i, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00143933. See

[Nel94d].

Nelder:1997:CDF

[Nel97] John Nelder. Contribution to discussion on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1): 108, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900124. See [Les97].

Nelder:1998:LEB

[Nel98a] J. A. Nelder. Letter to the Editor: A Bayesian analysis of regression models with continuous errors with application to longitudinal studies, by L. D. Broemeling and P. Cook, Statistics in Medicine, 16, 321–332 (1997). Statistics in Medicine, 17(2):241, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [BC97].

Nelder:1998:LEI

[Nel98b] J. A. Nelder. Letters to the Editors: The importance of marginality rules. Applied Statistics, 47(3):447–448, 1998. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0035-9254&src=ard&aid=122&iid=3&vid=47.

Nelder:1998:LCM

[Nel98c] John A. Nelder. A large class of models derived from generalized linear models. Statistics in Medicine, 17(23):2747–2753, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1998:STR

[Nel98d] John A. Nelder. The selection of terms in response-surface models: How strong is the weak-heredity principle? The American Statistician, 52(4):315–318, November 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.amstat.org/publications/tas/abstracts_
98/nelder.html;   http://www.jstor.org/stable/2685433.

Nelder:1999:SSS

[Nel99] John A. Nelder. From statistics to statistical science. Journal of the Royal Statistical Society. Series D (The Statistician), 48(2): 257–269, ???? 1999. CODEN ???? ISSN 0039-0526 (print),

1467-9884 (electronic). URL http://www.jstor.org/stable/
2681191.

Nelder:2000:FMR

[Nel00a] J. A. Nelder. Functional marginality and response-surface fitting. Journal of Applied Statistics, 27(1):109–112, January 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (elec-

tronic). URL http://www.catchword.co.uk/cgi-bin/cgi?ini=
carfax&body=linker&reqidx=/catchword/carfax/13600532/
v27n1/s8/p109.

Nelder:2000:QLP

[Nel00b] J. A. Nelder. Quasi-likelihood and pseudo-likelihood are not the same thing. Journal of Applied Statistics, 27(8):1007–1011, November 1, 2000. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelder:2000:ACT

[Nel00c] John A. Nelder. The analysis of contingency tables with one factor as the response: Round two. Journal of the Royal Statistical Society. Series D (The Statistician), 49(3):383–388, ???? 2000. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2681063.

Nelder:2001:C

[Nel01a] J. A. Nelder. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 50(2):209–211, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2681095.

Nelder:2001:LES

[Nel01b] John A. Nelder. Letter to the Editor: Statistics in medical journals: some recent trends by Douglas G. Altman, Statistics in Medicine 2000; 19: 3275–3289. Statistics in Medicine, 20(14): 2205, July 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [Alt00].

Nelder:2004:F

[Nel04a] J. A. Nelder. Foreword. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Nelder:2004:SEH

[Nel04b] John A. Nelder. A statistical examination of the Hastings Rarities. In Adams et al. [ACHS04c], pages 215–234. ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Nelder:2005:SQ

[Nel05] John Nelder. Synthesis and quality. Significance (Oxford, England), 2(2):75–77, June 2005. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Nelder:1984:MRP

[NF84] J. A. Nelder and E. L. Frome. Models for rates with Poisson errors. Biometrics, 40(4):1159–1162, December 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2531169.

Nelder:1994:DSQ

[NGL^{+}94] J. A. Nelder, T. Greenfield, H. J. Lenz, C. Chatfield, D. A. Preece, C. E. Lunneborg, M. C. Jones, J. Gower, R. A. Stone, M. C. Fessey, S. J. W. Evans, T. Lewis, A. S. C. Ehrenberg, D. J. Finney, A. M. Herzberg, P. Lovie, A. D. Lovie, R. J. Mackay, R. W. Oldford, I. W. Molenaar, P. C. Obrien, H. Rouanet, T. M. F. Smith, J. Tukey, M. Wise, and J. Zighera. Deconstructing statistical questions — discussion. Journal of the Royal Statistical Society Series A — Statistics in Society, 157(??):338–356, ???? 1994. ISSN 0035-9238.

Nelder:1994:RRR

[NGMS94] J. A. Nelder, E. I. George, U. E. Makov, and A. F. M. Smith. Reader reaction: a re-analysis of the pump-failure data [with response]. Scandinavian Journal of Statistics. Theory and Applications, 21(2):187–191, June 1994. CODEN SJSADG. ISSN 0303-6898 (print), 1467-9469 (electronic). URL https://www.jstor.

org/stable/4616310.

Nelder:1992:ESP

[NGN92] J. A. Nelder, Sander Greenland, and R. G. Newcombe. Estimating standardized parameters from generalized linear models. Statistics in Medicine, 11(4):559–560, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1992:LE

[NKA92] J. A. Nelder, Stena Kettl, and Harold Ascher. Letters to the editors. Applied Statistics, 41(3):595–600, 1992. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nelder:1992:LQL

[NL92] J. A. Nelder and Y. Lee. Likelihood, quasi-likelihood and pseudolikelihood: some comparisons. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):273–284, 1992. CO-DEN JSTBAJ. ISSN 0035-9246. URL http://links.jstor.

org/sici?sici=0035-9246(1992)54:1<273:LQAPSC>2.0.CO;2-
Y&origin=MSN.

Nelder:1995:CAF

[NL95] John A. Nelder and Peter W. Lane. The computer analysis of factorial experiments: In memorian — Frank Yates. The American Statistician, 49(4):382–385, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684580.

Nelder:1997:EQL

[NL97] J. A. Nelder and Y. Lee. Extended quasi-likelihood and estimating equations. In Selected Proceedings of the Symposium on Estimating Functions (Athens, GA, 1996), volume 32 of IMS Lecture Notes Monogr. Ser., pages 139–148. Institute of Mathematical Statistics, Hayward, CA, USA, 1997. ISBN 0-940600-44-7.

Nelder:1998:JMM

[NLB^{+}98] John A. Nelder, Youngjo Lee, Bo Bergman, Anders Hynén, A. Freek Huele, and Jan Engel. Joint modeling of mean and dispersion. Technometrics, 40(2):168–175, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270676.

Nelder:2008:LEW

[NLSS08] John A. Nelder, Viviana B. Lencina, Julio M. Singer, and Edward J. Stanek III. Letter to the Editors: What is the mixedmodels controversy? (with response). International Statistical Review = Revue Internationale de Statistique, 76(1):134–135, April 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.jstor.org/stable/27919588.

Nelder:1965:ESM

[NM65a] J. Nelder and R. Mead. Errata: “A simplex method for function minimization [MR3363409]. The Computer Journal, 8(1):27, 1965. CODEN CMPJA6. ISSN 1460-2067.

Nelder:1965:SMF

[NM65b] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7(4):308–313, 1965. CODEN CMPJA6. ISSN 1460-2067.

Nelder:1987:EQL

[NP87a] J. A. Nelder and D. Pregibon. An extended quasi-likelihood function. Biometrika, 74(2):221–232, June 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2336136.

Nelder:1987:EQF

[NP87b] J. A. Nelder and D. Pregibon. An extended quasilikelihood function. Biometrika, 74(2):221–232, 1987. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

Nelder:1974:STR

[NPT^{+}74] J. A. Nelder, G. P. Patil, J. L. Teugels, T. Dalenius, Saw Swee-Hock, and J. S. Siegel. Statistical training and research/formation et recherche statistiques. International Statistical Review = Revue Internationale de Statistique, 42(1):79–87, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1402687.

Nelder:1997:LE

[NR97] J. A. Nelder and G. J. S. Ross. Letters to the editors. Applied Statistics, 46(2):281–286, 1997. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0035-9254&src=ard&aid=066&iid=2&vid=46.

Nelder:1991:GLM

[NRCC91] J. A. Nelder, D. Ruppert, N. Cressie, and R. J. Carroll. Generalized linear models for enzyme-kinetic data. Biometrics, 47(4): 1605–1615, December 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2532412.

Nelder:1974:LE

[NS74] J. A. Nelder and David A. Swick. Letters to the editors. Applied Statistics, 23(2):232–233, 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nazareth:2002:GLV

[NT02] Larry Nazareth and Paul Tseng. Gilding the lily: a variant of the Nelder–Mead algorithm based on golden-section search. Computational optimization and applications, 22(1):133–144, 2002. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Nelder:1972:GLM

[NW72] J. A. Nelder and R. W. M. Wedderburn. Generalized linear models. Journal of the Royal Statistical Society. Series A (General), 135(3):370–384, ???? 1972. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/2344614.

Nawaz:2015:JNH

[NZE^{+}15] M. Nawaz, A. Zeeshan, R. Ellahi, S. Abbasbandy, and Saman Rashidi. Joules and Newtonian heating effects on stagnation point flow over a stretching surface by means of genetic algorithm and Nelder–Mead method. International Journal of Numerical Methods for Heat & Fluid Flow, 25(3):665–684, 2015.

Olsson:1975:NMS

[ON75] Donald M. Olsson and Lloyd S. Nelson. The Nelder–Mead simplex procedure for function minimization. Technometrics, 17(1):45– 51, February 1975. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1267998.

Ouria:2009:ANM

[OT09] Ahad Ouria and Mohammad M. Toufigh. Application of Nelder– Mead simplex method for unconfined seepage problems. Applied Mathematical Modelling, 33(9):3589–3598, 2009. CODEN AM-MODL. ISSN 0307-904X (print), 1872-8480 (electronic).

Pawitan:2004:LPC

[Paw04] Y. Pawitan. Likelihood perspectives in the consensus and controversies of statistical modelling and inference. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Payne:2004:ADS

[Pay04] R. Payne. Algorithms, data structures and languages: the computational ingredients for innovative analysis. In Adams et al. [ACHS04c], pages 95–118. ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Puhan:2026:OPV

[PB26] Janez Puhan and Árpád Bűrmen. On the optimal parameter values of the Nelder–Mead simplex algorithm. Journal of Computational and Applied Mathematics, 484:Paper No. 117521, 18, 2026. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Price:2002:CVN

[PCB02] C. J. Price, I. D. Coope, and D. Byatt. A convergent variant of the Nelder–Mead algorithm. Journal of Optimization Theory and Applications, 113(1):5–19, 2002. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Peren:2024:NON

[Per24] Franz W. Peren. Nonlinear optimization: the Nelder–Mead simplex search procedure. In Operations research and management— quantitative methods for planning and decision-making in business and economics, Springer Texts Bus. Econ., pages 187–199. Springer, Cham, 2024. ISBN 3-031-47205-5; 3-031-47206-3.

Peren:2026:NON

[Per26] Franz W. Peren. Nonlinear optimization: the Nelder–Mead simplex search procedure. In Operations research and management— quantitative methods for planning and decision-making in business and economics, Springer Texts Bus. Econ., pages 187–199. Springer, Cham, 2026. ISBN 3-032-08077-0; 3-032-08078-9.

Plackett:1984:BRB

[Pla84] R. L. Plackett. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society.

Series A (General), 147(3):526–527, ???? 1984. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.jstor.org/stable/2981609.

Pandey:2023:PNI

[PP23] Vivek Pandey and Sudhir K. Pandey. PY-Nodes: an ab-initio Python code for searching nodes in a material using Nelder–Mead’s simplex approach. Computer Physics Communications, 283(??): ??, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465522002892.

Pregibon:1984:RPM

[Pre84] Daryl Pregibon. Review: P. McCullagh, J. A. Nelder, Generalized Linear Models. Annals of Statistics, 12(4):1589–1596, December 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176346819.

Price:2021:MNM

[Pri21] C. J. Price. A modified Nelder–Mead barrier method for constrained optimization. Numerical Algebra, Control and Optimization, 11(4):613–631, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http://aimsciences.org/

article/doi/10.3934/naco.2020058.

Ripley:2004:SAL

[Rip04] B. D. Ripley. Selecting amongst large classes of models. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Robinson:2007:BRBb

[Rob07] Andrew Robinson. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-Likelihood, by Y. Lee; J. A. Nelder; Y. Pawitan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(3):855–856, July 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/4623207.

Rodriguez:1995:CJN

[RTW95] Robert Rodriguez, Randall Tobias, and Russell Wolfinger. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):97–101, June 1995.

CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00143938. See [Nel94d].

Ren:2013:HED

[RWJ13] Aihong Ren, Yuping Wang, and Fei Jia. A hybrid estimation of distribution algorithm and Nelder–Mead simplex method for solving a class of nonlinear bilevel programming problems. Journal of Applied Mathematics, pages Art. ID 378568, 10, 2013. ISSN 1110-757X (print), 1687-0042 (electronic).

Schuenemeyer:1992:BRB

[Sch92] John H. Schuenemeyer. Book review: Generalized Linear Models by P. McCullagh; J. A. Nelder. Technometrics, 34(2):224, May 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1269238.

Searle:1995:CJN

[Sea95] Shayle R. Searle. Comments on J. A. Nelder, ‘The statistics of linear models: back to basics’. Statistics and Computing, 5(2):103– 107, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-

1375 (electronic). URL http://link.springer.com/article/
10.1007/BF00143939. See [Nel94d].

Suhail:2024:IDH

[SEABZ24] Enas Suhail, Mahmoud El-Alem, Omar Bazighifan, and Ahmed Zekri. Improving diversification by a hybrid bat-Nelder–Mead algorithm and DDE for rapid convergence to solve global optimization. AIMS Mathematics, 9(12):35655–35677, 2024. ISSN 2473-6988.

Selvan:2015:NIC

[Sel15] Suviseshamuthu Easter Selvan. Nonsmooth ICA contrast minimization using a Riemannian Nelder–Mead method. IEEE Transactions on Neural Networks and Learning Systems, 26(1):177–183, 2015.

Senn:2003:CJN

[Sen03] Stephen Senn. A conversation with John Nelder. Statistical Science, 18(1):118–131, February 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1056397489.

Senn:2004:JNG

[Sen04] S. Senn. John Nelder: From general balance to generalised models (both linear and hierarchical). In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (ebook). LCCN QA276.A1 M48 2004.

Senn:2019:JAN

[Sen19] Stephen Senn. John Ashworth Nelder. 8 October 1924–7 August 2010. Biographical Memoirs of Fellows of the Royal Society, 67:307–326, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0013.

Sarakhsi:2016:NHA

[SGK16] M. Khojaste Sarakhsi, S. M. T. Fatemi Ghomi, and B. Karimi. A new hybrid algorithm of scatter search and Nelder–Mead algorithms to optimize joint economic lot sizing problem. Journal of Computational and Applied Mathematics, 292(??):387–401, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042715003891.

Smith:2001:SSM

[SNB^{+}01] Andrew J. Smith, John Nelder, Andreas Buja, Zahid Malik, Lisa Tweedie, and Robert Spence. Sampling schemes for model visualization. Journal of Computational and Graphical Statistics, 10(3): 545–554, 2001. ISSN 1061-8600 (print), 1537-2715 (electronic).

Singer:2001:CAN

[SS01] Sanja Singer and Saša Singer. Complexity analysis of Nelder– Mead search iterations. In Applied mathematics and computation (Dubrovnik, 1999), pages 185–196. Univ. Zagreb Dep. Math., Zagreb, 2001. ISBN 953-6076-70-5.

Singer:2004:EIN

[SS04] Saša Singer and Sanja Singer. Efficient implementation of the Nelder–Mead search algorithm. ANACM. Applied Numerical Analysis and Computational Mathematics, 1(3):524–534, 2004. ISSN 1611-8170 (print), 1611-8189 (electronic).

Turksen:2013:EES

[TA13] Özlem Türkşen and Ayşen Apaydin. Estimating the earthquake source parameters: simulated annealing versus Nelder–Mead simplex algorithm. Communications de la Faculté des Sciences de l’Université d’Ankara. Séries A1. Mathematics and Statistics. Communications. Faculty of Sciences. University of Ankara, 62 (2):53–66, 2013. ISSN 1303-5991.

Tekile:2021:CEM

[TFB21] Hailemariam Abebe Tekile, Michele Fedrizzi, and Matteo Brunelli. Constrained eigenvalue minimization of incomplete pairwise comparison matrices by Nelder–Mead algorithm. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/222.

Tomick:1995:CNM

[Tom95] John James Tomick. On convergence of the Nelder–Mead simplex algorithm for unconstrained stochastic optimization. Thesis (Ph.D.), The Pennsylvania State University, State College,

PA, USA, 1995. 188 pp. URL https://www.proquest.com/
pqdtglobal/docview/304210989.

Takenaga:2023:PIN

[TOO23] Shintaro Takenaga, Yoshihiko Ozaki, and Masaki Onishi. Practical initialization of the Nelder–Mead method for computationally expensive optimization problems. Optimization Letters, 17(2): 283–297, 2023. ISSN 1862-4472 (print), 1862-4480 (electronic).

Ulin:1958:CNM

[Uli58] Bengt Ulin. On a conjecture of Nelder in mathematical statistics. Compositio Mathematica, 13:148–149, 1958. CODEN CMPMAF. ISSN 0010-437X (print), 1570-5846 (electronic).

VanEeuwijk:1995:TDS

[Van95] F. A. Van Eeuwijk. On the tenability of distinctions: some comments on ‘The statistics of linear models: back to basics’ by John Nelder. Statistics and Computing, 5(2):93–95, June 1995. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF00143937. See [Nel94d].

Wakefield:2004:NLR

[Wak04] J. C. Wakefield. Non-linear regression modelling and inference. In Adams et al. [ACHS04c], page ?? ISBN 1-281-86647-4, 1-86094-463-9, 1-86094-541-4 (e-book). LCCN QA276.A1 M48 2004.

Watson:1986:BRG

[Wat86] G. S. Watson. Book review: Generalized Linear Models (P. Mccullagh and J. A. Nelder). SIAM Review, 28(1):128–130, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wessing:2019:PIC

[Wes19] Simon Wessing. Proper initialization is crucial for the Nelder– Mead simplex search. Optimization Letters, 13(4):847–856, 2019. ISSN 1862-4472 (print), 1862-4480 (electronic).

Williams:1984:BRG

[Wil84] D. A. Williams. Book review: Generalized Linear Models, by P. McCullagh, J. A. Nelder. Biometrics, 40(2):566, June 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2531415.

Wirjanto:1997:CFD

[Wir97] T. S. Wirjanto. Comment on A. F. Desmond’s paper: Optimal estimating functions, quasi-likelihood and statistical modelling. Journal of Statistical Planning and Inference, 60(1):109– 113, April 30, 1997. CODEN JSPIDN. ISSN 0378-3758 (print),

1873-1171 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0378375897900136. See [Des97a].

Wright:2012:NMO

[Wri12] Margaret H. Wright. Nelder, Mead, and the other simplex method. Documenta mathematica, pages 271–276, 2012. ISSN 1431-0643.

Wyers:2013:BDN

[WSKF13] Eric J. Wyers, Michael B. Steer, C. T. Kelley, and Paul D. Franzon. A bounded and discretized Nelder–Mead algorithm suitable for RFIC calibration. IEEE Transactions on Circuits and Systems I: Regular Papers, 60(7):1787–1799, 2013. ISSN 1549-8328 (print), 1558-0806 (electronic).

Wang:2015:FBS

[WYWW15] Ze Wang, Limin Yang, Chi Man Wong, and Feng Wan. Fast basis searching method of adaptive Fourier decomposition based on Nelder–Mead algorithm for ECG signals. In Advances in neural networks—ISNN 2015, volume 9377 of Lecture Notes in Comput. Sci., pages 305–314. Springer, Cham, 2015. ISBN 3-319-25393-X; 3-319-25392-1.

Xiao:2014:MDB

[XD14] Hong Feng Xiao and Ji An Duan. Multi-direction-based Nelder– Mead method. Optimization, 63(7):1005–1026, 2014. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898.

Yang:2024:JIJ

[YCH24] Yuxin Yang, Xiaoming Chen, and Yinhe Han. JANM-IK: Jacobian argumented Nelder–Mead algorithm for inverse kinematics and its hardware acceleration. IEEE Computer Architecture Letters, 23 (1):45–48, January/June 2024. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Young:1976:LE

[YYB^{+}76] Kan Hua Young, Lih Ying Young, A. E. Brandt, C. E. Werts, D. A. Rock, R. L. Linn, K. C. Joreskog, R. W. Farebrother, Oliver D. Anderson, Federico J. O’Reilly, J. A. Nelder, Nathan Mantel, Mark J. Nicolich, and John Neter. Letters to the Editor. The American Statistician, 30(2):101–104, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683805.

Zhang:2007:SMU

[ZHZ^{+}07] Kai Zhang, Ming You Huang, Ran Zhang, Gang Wang, and Jin Yue Gao. Simulation measurement of ultrafine nanoparticle size: Nelder–Mead simplex method. Gongcheng Shuxue Xuebao, 24(6):977–986, 2007. ISSN 1005-3085.

Zhu:2022:IHH

[ZZP^{+}22] Cheng Zhu, Yong Zhang, Xuhua Pan, Qi Chen, and Qingyu Fu. Improved Harris hawks optimization algorithm based on quantum correction and Nelder–Mead simplex method. Mathematical Biosciences and Engineering. MBE, 19(8):7606–7648, 2022.